For a polynomial of the form ax2 + bx + c = 0 to have real solutions, √(b2 – 4ac) must be a real number. Which of the following polynomials have real solutions?
a. x2 – 12x + 3 = 0

Answer by stanbon(75887)

x^2 – 12x + 3 = 0
-----------------
Checking the discriminant you get:
b^2-4ac = 12^2-4*1*3 = 132
------------
Since the discriminant is greater than zero the quadratic
equation has two unequal Real solutions.
=================
Cheers,
Stan H.